Na2SO4 * xH2O, the hydrated salt is overheated and the anhydrous salt thermally decompose, one product being a gas; will the reported value of "x" be too high, too low on the unaffected? explain
1
Expert's answer
2010-10-13T12:50:33-0400
Na2SO4*xH2O = (dehydratation) = Na2SO4 + xH2O.x = n(H2O)/n(Na2SO4*xH2O), where n(H2O) and n(Na2SO4*xH2O) the number of moles of the water and of the hydrated salt respectively.The sodium sulfate exists mostly in the dehydrate form: Na2SO4*10H2O, so x would be equal to 10.